OpenStudy (anonymous):

If you flip a coin 7 times, what is the probability of it being 4 tails and 3 heads? -Can this be done without binomial?

OpenStudy (anonymous):

you mean without \(\binom{7}{3}\)?

OpenStudy (anonymous):

you have \(2^7=128\) elements in your sample space, so i would not want to count all the possibilities by hand.

OpenStudy (anonymous):

Like, the formula nCx p^x q^(n-x) can be used but could you also do: P( Four Tails) x P(3 heads) ?

OpenStudy (anonymous):

answer is easy enough to get since 128 is your denominator and \[\dbinom{7}{3}=\frac{7\times 6\times 5}{3\times 2}=35\]

OpenStudy (anonymous):

you don't need all that since \(p=1-p=\frac{1}{2}\)

OpenStudy (kinggeorge):

I'm pretty sure the binomial distribution would at least be the easiest way to do it.

OpenStudy (anonymous):

and no, it is not the product

OpenStudy (anonymous):

why is that?

OpenStudy (anonymous):

you have 128 elements in your sample space all equally likely, so i would not even think of the formula that you wrote. it is just \(\frac{35}{128}\)

OpenStudy (anonymous):

first off, i am not even sure what \(P(\text{4 tails})\) means. four tails in how many tries?

OpenStudy (anonymous):

How did you get this?

OpenStudy (anonymous):

saying" exactly 4 tails in 7 tries" is the same as saying "4 tails and 3 heads".

OpenStudy (anonymous):

how did i get what?

Directrix (directrix):

Pascal's Triangle would give the 35 factor.

OpenStudy (kinggeorge):

I'd like to interject once again saying that what what satellite is doing works very well in this case. However, if you have a biased coin, the binomial distribution is by far and away the best way to do it. (To my knowledge)

OpenStudy (anonymous):

yes sorry i am trying to understand how satellite73 came to the anwer though. I cannot see it easily. I'm quite slow

OpenStudy (anonymous):

yes of course. and you are uisng it in this case too, but it is not necessary to write \((\frac{1}{2})^4\times (1-\frac{1}{2})^3\) when it is simply \((\frac{1}{2})^7\)

OpenStudy (anonymous):

@bobobobobb there are 35 possible cominations of 4 heads and 3 tails in 7 tosses. since each outcome is equally likely, and there are 128 outcomes altogether, your answer is \(\frac{35}{128}\)

OpenStudy (anonymous):

How did you get 35?

OpenStudy (kinggeorge):

35=7C3

OpenStudy (anonymous):

oh thank you @KingGeorge

OpenStudy (anonymous):

and thanks @satellite73

OpenStudy (anonymous):

yw 35 was computed here \[\dbinom{7}{3}=\frac{7\times 6\times 5}{3\times 2}=35\]